# Localized RETE for Incremental Graph Queries with Nested Graph Conditions

### v1

@@ 1. Introduction
This article is an extension of our paper at ICGT 2024 [BG24a]. The main increment of this article compared to the conference paper version is the approach for localizing execution of graph queries with nested graph conditions presented in Section 4, which is also covered by an extended evaluation in Section 5. Besides small improvements to text and figures, we have also added a theorem for recognizing locally relevant query results as another minor extension in Section 3.
@@ 2.1. Graphs and Graph Queries
We say that $G$ is edge-dominated if $\forall e_{TG}\in E^{TG}:|\{e\in E^{G}|type(e)=e_{TG}\}|\geq max(|\{v\in V^{G}|type(v)=s^{TG}(e_{TG})\}|,|\{v\in V^{G}|type(v)=t^{TG}(e_{TG})\}|)$.
A model is then simply given by a (typed) graph. In the remainder of the article, we therefore use the terms graph and model interchangably.
A plain graph query as considered in this article is characterized by a (typed) query graph $Q$ and can be executed over a (typed) host graph $H$ by finding (typed) graph morphisms $m:Q\rightarrow H$. We also call these morphisms matches and denote the set of all matches from $Q$ into $H$ by $\mathcal{M}^{Q}_{H}$. Typically, a set of matches is considered a complete query result if it contains all matches in $\mathcal{M}^{Q}_{H}$:
@@
(Completeness of Plain Query Results) We say that a set of matches $M$ from query graph $Q$ into host graph $H$ is complete iff $M=\mathcal{M}^{Q}_{H}$.
In some cases, the formalism of plain graph queries may not be sufficiently expressive to specify the desired query, for instance if matches for the query should not be extensible by a specific negative pattern. To increase expressiveness, a plain graph query $Q$ can therefore be equipped with a nested graph condition [HP09] $\psi$, yielding the extended graph query $(Q,\psi)$.
@@
  • $\exists(a:Q\rightarrow Q^{\prime},\psi^{\prime})$, where $a:Q\rightarrow Q^{\prime}$ is a (typed) graph monomorphism, is a nested graph condition that is satisfied iff there exists a (typed) graph morphism $m^{\prime}:Q^{\prime}\rightarrow H$ such that $m=m^{\prime}\circ a$ and $m^{\prime}$ satisfies the nested graph condition $\psi^{\prime}$, that is, $m\models\exists(a:Q\rightarrow Q^{\prime},\psi^{\prime})\Leftrightarrow(\exists m^{\prime}\in\mathcal{M}^{Q^{\prime}}_{H}:m^{\prime}\models\psi^{\prime}\wedge m=m^{\prime}\circ a)$.
Equipping the query graph $Q$ with the nested graph condition $\psi$ restricts the set of valid matches for the extended graph query $(Q,\psi)$ to those matches for $Q$ that also satisfy $\psi$. Correctness of query results for extended graph queries can hence be defined as follows:
@@ 2.2. Incremental Graph Queries with RETE
We say that $\mathcal{C}$ is consistent for a RETE node $n\in V^{N}$ and host graph $H$ iff $\mathcal{C}(n)=\mathcal{R}(n,N,H,\mathcal{C})$. If $\mathcal{C}$ is consistent for all $n\in V^{N}$, we say that $\mathcal{C}$ is consistent for $H$. If $H$ is clear from the context, we simply say that $\mathcal{C}$ is consistent.
@@ 3. Incremental Plain Queries over Subgraphs
(Completeness of Plain Query Results under Subgraphs) We say that a set of matches $M$ from query graph $Q$ into host graph $H$ is complete under a subgraph $H_{p}\subseteq H$ iff $\forall m\in\mathcal{M}^{Q}_{H}:m(Q)\cap H_{p}\neq\emptyset\Rightarrow m\in M$.
@@ 3.1. Marking-sensitive RETE
Therefore, we extend the standard RETE mechanism by a natural number marking for matches. Intuitively, we will use this marking to encode up to which height in a RETE join tree derived matches should trigger the retrieval of complementary elements and matches to guarantee completeness with respect to Definition 3. In our extension, an intermediate result is therefore characterized by a tuple $(m,\phi)$ of a match $m$ and a marking $\phi\in\overline{\mathbb{N}}$, where we define $\overline{\mathbb{N}}:=\mathbb{N}\cup\{\infty\}$. A marking-sensitive configuration is then given by a function $\mathcal{C}^{\Phi}:V^{N}\rightarrow\mathcal{P}(\mathcal{M}_{\Omega}\times\overline{\mathbb{N}})$.
@@
Formally, for a host graph $H$, a relevant subgraph $H_{p}\subseteq H$, a containing marking-sensitive RETE net $(N^{\Phi},p^{\Phi})$, and a configuration $\mathcal{C}^{\Phi}$ for $(N^{\Phi},p^{\Phi})$, we define the target result sets of marking-sensitive RETE nodes as follows:
@@
  • The target result set of a marking filter node $[\phi>i_{max}]^{\Phi}$ with marking-sensitive dependency $n^{\Phi}_{\alpha}$ is given by $\mathcal{R}^{\Phi}([\phi>i_{max}]^{\Phi},N^{\Phi},H,H_{p},\mathcal{C}^{\Phi})=\{(m,\phi_{\alpha})|(m,\phi_{\alpha})\in\mathcal{C}^{\Phi}(n^{\Phi}_{\alpha})\wedge\phi_{\alpha}>i_{max}\}$.
  • The target result set of a forward navigation node $[v\rightarrow_{n}w]^{\Phi}$ with $Q=query([v\rightarrow_{n}w]^{\Phi})=(\{v,w\},\{e\},\{(e,v)\},\{(e,w)\})$ and marking-sensitive dependency $n^{\Phi}_{v}$ with $Q_{v}=query(n^{\Phi}_{v})=(\{v\},\emptyset,\emptyset,\emptyset)$ is given by $\mathcal{R}^{\Phi}([v\rightarrow_{n}w]^{\Phi},N^{\Phi},H,H_{p},\mathcal{C}^{\Phi})=\{(m,\phi_{max})|m\in\mathcal{M}^{Q}_{H}\wedge\exists(m_{v},\phi_{max})\in\mathcal{C}^{\Phi}(n^{\Phi}_{v}):m_{v}=m|_{Q_{v}}\wedge\phi_{max}=max(\{\phi^{\prime}|(m_{v}^{\prime},\phi^{\prime})\in\mathcal{C}^{\Phi}(n^{\Phi}_{v})\wedge m_{v}^{\prime}=m|_{Q_{v}}\})\}\}$.
  • The target result set of a backward navigation node $[w\leftarrow_{n}v]^{\Phi}$ with $Q=query([w\leftarrow_{n}v]^{\Phi})=(\{v,w\},\{e\},\{(e,v)\},\{(e,w)\})$ and marking-sensitive dependency $n^{\Phi}_{w}$ with $Q_{w}=query(n^{\Phi}_{w})=(\{w\},\emptyset,\emptyset,\emptyset)$ is given by $\mathcal{R}^{\Phi}([w\leftarrow_{n}v]^{\Phi},N^{\Phi},H,H_{p},\mathcal{C}^{\Phi})=\{(m,\phi_{max})|m\in\mathcal{M}^{Q}_{H}\wedge\exists(m_{w},\phi_{max})\in\mathcal{C}^{\Phi}(n^{\Phi}_{w}):m_{w}=m|_{Q_{w}}\wedge\phi_{max}=max(\{\phi^{\prime}|(m_{w}^{\prime},\phi^{\prime})\in\mathcal{C}^{\Phi}(n^{\Phi}_{w})\wedge m_{w}^{\prime}=m|_{Q_{w}}\})\}\}$.
@@
To obtain query results in the format of the standard RETE approach, we define the stripped result set of a marking-sensitive RETE node $n^{\Phi}$ for a marking-sensitive configuration $\mathcal{C}^{\Phi}$ as $\mathcal{R}^{\Phi}_{X}(n^{\Phi},\mathcal{C}^{\Phi})=\{m|(m,\phi)\in\mathcal{C}^{\Phi}(n^{\Phi})\}$, that is, the set of matches that appear in tuples in the node’s current result set in $\mathcal{C}^{\Phi}$.
@@ 3.2. Localized Search with Marking-sensitive RETE
Importantly, the marking-sensitive vertex input nodes of the local navigation structure are executed over the relevant subgraph, whereas the forward and backward navigation nodes are executed over the full model. Intuitively, the local navigation structure thus takes the role of the edge input node, but initially only extracts edges that are adjacent to a vertex in the relevant subgraph.
@@
The result of applying $localize$ to the example RETE net in Figure 1 is displayed in Figure 3 (right). It consists of a marking sensitive join and two local navigation structures connected via request projection structures.
@@
A consistent configuration for a localized RETE net then indeed guarantees query results that are complete according to Definition 3:
@@
Let $H$ be a graph, $H_{p}\subseteq H$, $(N,p)$ a well-formed RETE net, and $Q=query(p)$. Furthermore, let $\mathcal{C}^{\Phi}$ be a consistent configuration for the localized RETE net $(N^{\Phi},p^{\Phi})=localize(N,p)$. The set of matches from $Q$ into $H$ given by the stripped result set $\mathcal{R}^{\Phi}_{X}(p^{\Phi},\mathcal{C}^{\Phi})$ is then complete under $H_{p}$.
@@
Let $H$ be a graph, $H_{p}\subseteq H$, $(N,p)$ a well-formed RETE net, and $Q=query(p)$. Furthermore, let $\mathcal{C}^{\Phi}$ be a consistent configuration for the localized RETE net $(N^{\Phi},p^{\Phi})=localize(N,p)$. It then holds that $\forall(m,\phi)\in\mathcal{C}^{\Phi}(p^{\Phi}):m(Q)\cap H_{p}\neq\emptyset\Leftrightarrow\phi=\infty$.
@@
Let $H$ be a graph, $H_{p}\subseteq H$, $(N,p)$ a well-formed RETE net, and $\mathcal{C}^{\Phi}_{0}$ an arbitrary starting configuration. Executing $(N^{\Phi},p^{\Phi})=localize(N,p)$ via $O=order(N^{\Phi})$ then yields a consistent configuration $\mathcal{C}^{\Phi}=execute(O,N^{\Phi},H,H_{p},\mathcal{C}^{\Phi}_{0})$.
@@
Let $H$ be a graph, $H_{p}\subseteq H$, $(N,p)$ a well-formed RETE net, and $Q=query(p)$. Furthermore, let $\mathcal{C}^{\Phi}_{0}$ be an arbitrary starting configuration for the marking-sensitive RETE net $(N^{\Phi},p^{\Phi})=localize(N,p)$ and $\mathcal{C}^{\Phi}=execute(order(N^{\Phi}),N^{\Phi},H,H_{p},\mathcal{C}^{\Phi}_{0})$. The set of matches from $Q$ into $H$ given by $\mathcal{R}^{\Phi}_{X}(p^{\Phi},\mathcal{C}^{\Phi})$ is then complete under $H_{p}$.

### v2

@@ 1. Introduction
This article is an extension of our paper at ICGT 2024 [BG24a]. The main increment of this article compared to the conference paper version is the approach for localizing execution of graph queries with nested graph conditions presented in Section 4, which is also covered by an extended evaluation in Section 5. Besides small improvements to text and figures and the inclusion of further examples, we have also added a theorem for recognizing locally relevant query results as another minor extension in Section 3.
@@ 2.1. Graphs and Graph Queries
A model is then simply given by a typed graph. In the remainder of the article, we therefore use the terms graph and model interchangably and implicitly assume typing when talking about graphs.
In practice, graphs often contain more edges than nodes, a characteristic that may influence the performance of certain algorithms. To capture this property for each individual edge type in a typed graph, we introduce the notion of edge-dominated graphs:
@@
(Edge-Dominated Typed Graphs) We say that a typed graph $G$ is edge-dominated if $\forall e_{TG}\in E^{TG}:|\{e\in E^{G}|type(e)=e_{TG}\}|\geq max(|\{v\in V^{G}|type(v)=s^{TG}(e_{TG})\}|,|\{v\in V^{G}|type(v)=t^{TG}(e_{TG})\}|)$.
The above definition will be relevant in the context of an analytical evaluation of our localized querying technique, which technically requires graphs to be edge-dominated in order to guarantee desirable characteristics regarding computational complexity.
A plain graph query as considered in this article is characterized by a query graph $Q$ and can be executed over a host graph $H$ by finding graph morphisms $m:Q\rightarrow H$. We also call these morphisms matches and denote the set of all matches from $Q$ into $H$ by $\mathcal{M}^{Q}_{H}$. Typically, a set of matches is considered a complete query result if it contains all matches in $\mathcal{M}^{Q}_{H}$:
{defi}
(Completeness of Plain Query Results) We say that a set of matches $M$ from query graph $Q$ into host graph $H$ is complete iff $\mathcal{M}^{Q}_{H}\subseteq M$.
Note that by virtue of $M$ being defined as a set of matches from $Q$ into $H$, the statement $\mathcal{M}^{Q}_{H}\subseteq M$ in the above definition implies that $M=\mathcal{M}^{Q}_{H}$.
In some cases, the formalism of plain graph queries may not be sufficiently expressive to specify the desired query, for instance if matches for the query should not be extensible by a specific negative pattern. To increase expressiveness, a plain graph query $Q$ can therefore be equipped with a nested graph condition $\psi$ [HP09], yielding the extended graph query $(Q,\psi)$.
@@
  • $\exists(a:Q\rightarrow Q^{\prime},\psi^{\prime})$, where $a:Q\rightarrow Q^{\prime}$ is a graph monomorphism, is a nested graph condition that is satisfied iff there exists a graph morphism $m^{\prime}:Q^{\prime}\rightarrow H$ such that $m=m^{\prime}\circ a$ and $m^{\prime}$ satisfies the nested graph condition $\psi^{\prime}$, that is, $m\models\exists(a:Q\rightarrow Q^{\prime},\psi^{\prime})\Leftrightarrow(\exists m^{\prime}\in\mathcal{M}^{Q^{\prime}}_{H}:m^{\prime}\models\psi^{\prime}\wedge m=m^{\prime}\circ a)$.
As in [BG23], we also consider the case where the morphism $a$ of a nested graph condition $\exists(a:Q\rightarrow Q^{\prime},\psi^{\prime})$ can be a partial graph morphism from some subgraph $Q_{p}\subseteq Q$ of $Q$ into $Q^{\prime}$. In this case, satisfaction of the nested graph condition is given by $m\models\exists(a:Q\rightarrow Q^{\prime},\psi^{\prime})\Leftrightarrow(\exists m^{\prime}\in\mathcal{M}^{Q^{\prime}}_{H}:m^{\prime}\models\psi^{\prime}\wedge m|_{Q_{p}}=m^{\prime}\circ a)$.
Equipping the query graph $Q$ with the nested graph condition $\psi$ restricts the set of valid matches for the extended graph query $(Q,\psi)$ to those matches for $Q$ that also satisfy $\psi$. In contrast to plain graph queries, completeness alone is thus insufficient as a criterion for describing what constitutes a desirable query result. Specifically, when talking about extended graph queries, a query result should not only contain all admissible matches, but also no matches that violate the equipped nested graph condition. In this article, we will hence talk about completeness with respect to the results of plain graph queries and correctness in the case of extended graph queries, with the latter defined as follows:
@@ 2.2. Incremental Graph Queries with RETE
We say that $\mathcal{C}$ is consistent for a RETE node $n\in V^{N}$ and host graph $H$ iff $\mathcal{C}(n)=\mathcal{R}(n,N,H,\mathcal{C})$. We furthermore define a consistent configuration as a configuration that is consistent for all nodes of the RETE net:
{defi}
(Consistent RETE Net Configurations) We say that a configuration $\mathcal{C}:V^{N}\rightarrow\mathcal{P}(\mathcal{M}_{\Omega})$ for a RETE net $(N,p)$ is consistent for a host graph $H$ iff $\forall n\in V^{N}:\mathcal{C}(n)=\mathcal{R}(n,N,H,\mathcal{C})$.
If $H$ is clear from the context, we simply say that $\mathcal{C}$ is consistent.
@@
Figure 3 shows an example extended graph query on the left. The extended query searches for the same pattern as the plain query in Figure 1, but also requires the class node to be connected to an interface node, without this additional node being part of the main match. Therefore, the query is defined as (Q, $\exists(a:Q\rightarrow Q^{\prime},\texttt{true}))$ with a partial graph morphism $a$ that maps only the class node in $Q$ to the class node in $Q^{\prime}$.
The corresponding RETE net is visualized on the right of Figure 3. For computing matches for the base pattern, the RETE net replicates the structure in Figure 1. The nested graph condition is implemented via a search for the pattern $Q^{\prime}$, which consists of a single edge from the node $c$ to $i$, and a semi-join node at the top of the net, which checks the matches for $Q$ against the matches for $Q^{\prime}$.
p:Pkgc:Classf:Fieldi:Intfcefeie Q Q
[…]
Figure 3. Example extended graph query $(Q,\exists(a:Q\rightarrow Q^{\prime},\texttt{true}))$ (left) and corresponding RETE net (right)
@@ 3. Incremental Plain Queries over Subgraphs
(Completeness of Plain Query Results under Subgraphs) We say that a set of matches $M$ from query graph $Q$ into host graph $H$ is complete under a subgraph $H_{p}\subseteq H$ iff $\{m\in\mathcal{M}^{Q}_{H}|m(Q)\cap H_{p}\neq\emptyset\}\subseteq M$.
@@ 3.1. Marking-sensitive RETE
This approach is based on the observation that, in order for a join node to create a match that involves elements from the relevant subgraph, at least one of the constituting matches already has to involve at least one such element. Intuitively, our localized querying technique then works as follows: At the leaves of a RETE join tree for some plain graph query, we initially compute only matches that touch the relevant subgraph via local edge navigation. We then also implement a mechanism that enables a lazy computation of certain complementary partial matches that do not touch the relevant subgraph at inner nodes of the join tree. This ensures completeness of results with respect to Definition 3.
[…]
Moreover, such an approach would create cyclic requirement relationships between the matches $m_{l}$ and $m_{r}$, where the presence of $m_{l}$ in the result for $n_{l}$ mandates the presence of $m_{r}$ in the result for $n_{r}$ and vice-versa. This may prevent the unloading of matches when elements are removed from the relevant subgraph.
To mitigate these issues, we extend the standard RETE mechanism by a marking for matches in the form of a natural number. Intuitively, we will use this marking to encode up to which height in a RETE join tree derived matches should trigger the retrieval of complementary matches. An appropriate assignment of such markings then prevents the computation of the aforementioned superfluos matches as well as cyclic requirement relationships between matches, while still guaranteeing completeness according to Definition 3.
In our extension, an intermediate result in a marking-sensitive RETE net is therefore characterized by a tuple $(m,\phi)$ of a match $m$ and a marking $\phi\in\overline{\mathbb{N}}$, where we define $\overline{\mathbb{N}}:=\mathbb{N}\cup\{\infty\}$. A configuration for a marking-sensitive RETE net $(N^{\Phi},p^{\Phi})$ is then given by a function $\mathcal{C}^{\Phi}:V^{N^{\Phi}}\rightarrow\mathcal{P}(\mathcal{M}_{\Omega}\times\overline{\mathbb{N}})$.
@@
Consistency of marking-sensitive configurations is then defined analogously to the standard case:
{defi}
(Consistent Marking-Sensitive RETE Net Configurations) We say that a configuration $\mathcal{C}^{\Phi}:V^{N^{\Phi}}\rightarrow\mathcal{P}(\mathcal{M}_{\Omega}\times\overline{\mathbb{N}})$ for a marking-sensitive RETE net $(N^{\Phi},p^{\Phi})$ is consistent for a host graph $H$ with relevant subgraph $H_{p}\subseteq H$ iff $\forall n^{\Phi}\in V^{N^{\Phi}}:\mathcal{C}^{\Phi}(n^{\Phi})=\mathcal{R}^{\Phi}(n^{\Phi},N^{\Phi},H,H_{p},\mathcal{C}^{\Phi})$.
Note that technically, the definitions of marking-sensitive configurations and target result sets of marking-sensitive RETE nodes allow the same match to be associated with multiple markings for the same RETE node. However, as shown in [BG24b], the marking of a match in the target result set of a marking-sensitive RETE node is always unique in the constructions presented in this article.
@@
Formally, we define the target result sets of marking-sensitive RETE nodes as follows:
{defi}
(Target Result Sets of Marking-Sensitive RETE Nodes) Let $H$ be a host graph with relevant subgraph $H_{p}\subseteq H$, $(N^{\Phi},p^{\Phi})$ a containing marking-sensitive RETE net, and $\mathcal{C}^{\Phi}$ a configuration for $(N^{\Phi},p^{\Phi})$.
@@
  • The target result set of a marking filter node $[\phi>i_{min}]^{\Phi}$ with marking-sensitive dependency $n^{\Phi}_{\alpha}$ is given by $\mathcal{R}^{\Phi}([\phi>i_{min}]^{\Phi},N^{\Phi},H,H_{p},\mathcal{C}^{\Phi})=\{(m,\phi_{\alpha})|(m,\phi_{\alpha})\in\mathcal{C}^{\Phi}(n^{\Phi}_{\alpha})\wedge\phi_{\alpha}>i_{min}\}$.
  • The target result set of a forward navigation node $[v\rightarrow_{n}w]^{\Phi}$ with $Q=query([v\rightarrow_{n}w]^{\Phi})=(\{v,w\},\{e\},\{(e,v)\},\{(e,w)\})$ and marking-sensitive dependency $n^{\Phi}_{v}$ with $Q_{v}=query(n^{\Phi}_{v})=(\{v\},\emptyset,\emptyset,\emptyset)$ is given by $\mathcal{R}^{\Phi}([v\rightarrow_{n}w]^{\Phi},N^{\Phi},H,H_{p},\mathcal{C}^{\Phi})=\{(m,\phi)|m\in\mathcal{M}^{Q}_{H}\wedge(m|_{Q_{v}},\phi)\in\mathcal{C}^{\Phi}(n^{\Phi}_{v})\}$.
  • The target result set of a backward navigation node $[w\leftarrow_{n}v]^{\Phi}$ with $Q=query([w\leftarrow_{n}v]^{\Phi})=(\{v,w\},\{e\},\{(e,v)\},\{(e,w)\})$ and marking-sensitive dependency $n^{\Phi}_{w}$ with $Q_{w}=query(n^{\Phi}_{w})=(\{w\},\emptyset,\emptyset,\emptyset)$ is given by $\mathcal{R}^{\Phi}([w\leftarrow_{n}v]^{\Phi},N^{\Phi},H,H_{p},\mathcal{C}^{\Phi})=\{(m,\phi)|m\in\mathcal{M}^{Q}_{H}\wedge(m|_{Q_{w}},\phi)\in\mathcal{C}^{\Phi}(n^{\Phi}_{w})\}$.
@@
To obtain query results in the format of the standard RETE approach, we define the stripped result set of a marking-sensitive RETE node $n^{\Phi}$ for a marking-sensitive configuration $\mathcal{C}^{\Phi}$ as the set of matches that appear in tuples in the node’s current result set in $\mathcal{C}^{\Phi}$.
{defi}
(Stripped Result Sets of Marking-Sensitive RETE Nodes) The stripped result set of a marking-sensitive RETE node $n^{\Phi}$ for a marking-sensitive configuration $\mathcal{C}^{\Phi}$ is given by $\mathcal{R}^{\Phi}_{X}(n^{\Phi},\mathcal{C}^{\Phi})=\{m|(m,\phi)\in\mathcal{C}^{\Phi}(n^{\Phi})\}$.
@@ 3.2. Localized Search with Marking-sensitive RETE
Importantly, the marking-sensitive vertex input nodes of the local navigation structure are executed over the relevant subgraph, whereas the forward and backward navigation nodes are executed over the full model. Intuitively, the local navigation structure thus takes the role of the edge input node, but initially only extracts edges that are adjacent to a vertex in the relevant subgraph. Connecting additional dependencies to the union nodes $[\cup]^{\Phi}_{v}$ and $[\cup]^{\Phi}_{w}$ in further constructions then allows the extraction of additional edges that may be required to compute complete query results.
@@
The result of applying $localize$ to a RETE net consisting of a single join is displayed in Figure 4 (right). It consists of a marking sensitive join and two local navigation structures connected via request projection structures.
@@
[…]
Figure 5. Localized RETE net for the example query in Figure 1 with a consistent configuration for the host graph in Figure 6
The localized RETE net for the example query in Figure 1 is shown in Figure 5. Alongside the RETE net, the figure visualizes a consistent configuration for the host graph in Figure 6. In the example host graph, the relevant subgraph consists of the node $p_{1}$, which is indicated by a bold border.
With the relevant subgraph consisting of only a single node, the only intermediate result in any of the net’s vertex input nodes is the tuple $(m_{1.1.1},\infty)$ for the node $[p]^{\Phi}$. This intermediate result propagates up through the containing local navigation structure on the left, where it triggers the extraction of the match $m_{1.1}$ by the forward navigation node $[p\rightarrow_{n}c]^{\Phi}$. Via the connected request projection structure, the restricted match $m_{1.2.1}$ that consists only of the target node $c_{1}$ of the extracted edge is then propagated into the local navigation structure on the right. As a result, the match $m_{1.2}$ is extracted by the forward navigation node $[c\rightarrow_{n}f]^{\Phi}$, which is combined with the match $m_{1.1}$ by the join node $[\bowtie]^{\Phi}$ to form the complete match $m_{1}$.
Notably, the host graph contains a second match $m_{2}$ for the query pattern, which however does not touch the relevant subgraph. In this example, localization allows the RETE net to completely ignore this second match and skip any related computations.
Moreover, the assignment of the marking $1$ to the match $m_{1.2.1}$ in the request projection structure on the right prevents a propagation of any of the derived intermediate results back into the left local navigation structure. Thereby, potentially problematic cyclic requirement relationships of intermediate results are avoided.
As an example, consider the case where the node $p_{1}$ is removed from the relevant subgraph and this change is to be reflected in an updated configuration. Without the additional control of propagation provided by the marking filter and marking assignment nodes, the removal would not mandate an unloading of all associated intermediate results. Specifically, the result set for the union node below the backward navigation node $[c\leftarrow_{n}f]^{\Phi}$ would still contain the match $m_{1.2.1}$ and the backward navigation node would hence extract the match $m_{1.1}$ again, leading to the pollution of the RETE net’s configuration with superfluous intermediate results.
r1:Projp1:Pkgc1:Classf1:Fieldpececefep2:Pkgc2:Classf2:Fieldpecefe m1 m1.1 m1.1.1 m1.2 m1.2.1 m2
Figure 6. Example host graph with relevant subgraph marked in bold
Formally, a consistent configuration for a localized RETE net then indeed guarantees query results that are complete according to Definition 3:
@@
(Consistent configurations for RETE nets localized via $localize$ yield complete query results under the relevant subgraph.) Let $H$ be a graph, $H_{p}\subseteq H$, $(N,p)$ a well-formed RETE net, and $Q=query(p)$. Furthermore, let $\mathcal{C}^{\Phi}$ be a consistent configuration for the localized RETE net $(N^{\Phi},p^{\Phi})=localize(N,p)$. The set of matches from $Q$ into $H$ given by the stripped result set $\mathcal{R}^{\Phi}_{X}(p^{\Phi},\mathcal{C}^{\Phi})$ is then complete under $H_{p}$.
@@
(Matches are marked $\infty$ in query results of RETE nets localized via $localize$ iff they touch the relevant subgraph.) Let $H$ be a graph, $H_{p}\subseteq H$, $(N,p)$ a well-formed RETE net, and $Q=query(p)$. Furthermore, let $\mathcal{C}^{\Phi}$ be a consistent configuration for the localized RETE net $(N^{\Phi},p^{\Phi})=localize(N,p)$. It then holds that $\forall(m,\phi)\in\mathcal{C}^{\Phi}(p^{\Phi}):m(Q)\cap H_{p}\neq\emptyset\Leftrightarrow\phi=\infty$.
@@
(Execution of localized RETE nets via $order$ yields consistent configurations.) Let $H$ be a graph, $H_{p}\subseteq H$, $(N,p)$ a well-formed RETE net, and $\mathcal{C}^{\Phi}_{0}$ an arbitrary starting configuration. Executing $(N^{\Phi},p^{\Phi})=localize(N,p)$ via $O=order(N^{\Phi})$ then yields a consistent configuration $\mathcal{C}^{\Phi}=execute(O,N^{\Phi},H,H_{p},\mathcal{C}^{\Phi}_{0})$.
@@
(Execution of RETE nets localized via $localize$ yields complete query results under the relevant subgraph.) Let $H$ be a graph, $H_{p}\subseteq H$, $(N,p)$ a well-formed RETE net, and $Q=query(p)$. Furthermore, let $\mathcal{C}^{\Phi}_{0}$ be an arbitrary starting configuration for the marking-sensitive RETE net $(N^{\Phi},p^{\Phi})=localize(N,p)$ and $\mathcal{C}^{\Phi}=execute(order(N^{\Phi}),N^{\Phi},H,H_{p},\mathcal{C}^{\Phi}_{0})$. The set of matches from $Q$ into $H$ given by $\mathcal{R}^{\Phi}_{X}(p^{\Phi},\mathcal{C}^{\Phi})$ is then complete under $H_{p}$.